# MoDA: Multi-modal Diffusion Architecture for Talking Head Generation

### v1

@@ Abstract
Talking head generation with arbitrary identities and speech audio remains a crucial problem in the realm of digital humans and the virtual metaverse. Recently, diffusion models have become a popular generative technique in this field with their strong generation and generalization capabilities. However, several challenges remain for diffusion-based methods: 1) inefficient inference and visual artifacts, which arise from the implicit latent space of Variational Auto-Encoders (VAE), complicating the diffusion process; 2) authentic facial expressions and head movements, resulting from insufficient multi-modal information interaction. In this paper, MoDA handle these challenges by 1) defines a joint parameter space to bridge motion generation and neural rendering, and leverages flow matching to simplify the diffusion learning process; 2) introduces a multi-modal diffusion architecture to model the interaction among noisy motion, audio, and auxiliary conditions, ultimately enhancing overall facial expressiveness. Subsequently, a coarse-to-fine fusion strategy is adopted to progressively integrate different modalities, ensuring effective integration across feature spaces. Experimental results demonstrate that MoDA significantly improves video diversity, realism, and efficiency, making it suitable for real-world applications.
@@ 1 Introduction
Talking head generation is crucial for the metaverse and a wide range of applications, including immersive telepresence and augmented or virtual reality. This task involves generating a speaking portrait from a single image, guided by the corresponding audio and other auxiliary modalities. Lip synchronization for portrait generation has been extensively studied in computer vision. However, generating realistic and efficient talking head videos remains a persistent challenge.
@@
The early methods [24, 27] based on Generative Adversarial Networks (GANs) [11] were the primary solution for audio-to-portrait translation, demonstrating relatively accurate lip synchronization. However, these methods remain limited in capturing diverse speaking styles, expressive emotions, and natural head movements. Recently, diffusion models [7, 14, 33] have significantly advanced generative modeling, enabling the creation of highly diverse and high-fidelity images and videos. Motivated by this progress, significant research efforts [5, 3, 17, 31, 38, 43, 44, 2, 19] have been made aimed at extending these innovations to the development of portrait image animation.
@@
These methods are typically categorized into single-stage and two-stage frameworks. The former [5, 3, 17, 31, 38, 43] rely on a VAE-style latent space to directly generate video content from the audio input, integrating both motion and appearance in one pass. Despite being conceptually simple and trainable from end to end, two major limitations still persist, as shown in Fig. 2: 1) inefficient inference process. Most diffusion-based approaches must account for not only the diverse mapping of lip movements but also the rendering of visual appearances. This dual consideration significantly increases computational costs, making these methods impractical for real-time applications. 2) Limited capability for long-term video generation and visual artifacts. Methods like EMO [38] and Hallo [43] generate new video content by conditioning on previous frames. These prior frames provide motion information for subsequent generation and simultaneously leak appearance details, leading to the progressive accumulation of pixel errors and deterioration in overall video quality.
The two-stage methods [44, 2, 19] use diffusion models to generate low-dimensional intermediate motion representations, simplifying the learning process. However, the effectiveness of these methods is often compromised by inaccurate representations. This limitation mainly arises from the inherently one-to-many nature of the audio-to-motion task. The commonly used unidirectional cross-attention mechanism fails to support sufficient multi-modal interaction, leading to inconsistent motion sequences when conditioned on arbitrary identities and audio inputs, as discussed in Sec. 4.3.1. Motivated by recent lip-to-speech tasks [40, 26], visual information can provide complementary contextual cues to audio. Sufficient multi-modal interaction helps disambiguate motion generation, resulting in more accurate and identity-consistent outputs.
@@
In this paper, we propose a novel framework named MoDA, designed to generate realistic talking head videos from a single portrait image. MoDA first establishes a joint parameter space that bridges motion generation and neural rendering, consisting of a disentangled motion-appearance representation [13] along with the corresponding audio, emotion, and identity information. As shown in Fig. 3, MoDA introduces Multi-modal Diffusion Transformer (MMDiT) [9], equipped with rectified flow [21], as the backbone diffusion framework to facilitate multi-modal interaction. In this design, MoDA can dynamically adapt audio features based on motion, identity, and emotion, thereby improving the accuracy of motion generation.
@@
However, there are semantic overlaps between modalities within the joint parameter space: for example, audio inherently contains emotional and identity-related cues. Independently learned weights hinder the performance of the model. To facilitate effective feature fusion, a coarse-to-fine fusion architecture is further proposed to progressively enhance feature integration across modalities. Initially, the model maintains modality-specific weights to capture the unique characteristics of each condition. In the intermediate stage, audio and auxiliary conditions, such as identity and emotion, begin to share attention weights. Finally, all modalities are merged into a unified representation space, allowing holistic fusion and improving the overall generative expressiveness. Moreover, to further encourage motion diversity while maintaining temporal consistency, this paper provides an optional Adaptive Lip-motion Synchronization Expert (ALSE), which can be integrated during training.
The contributions of our method can be summarized as follows:
@@
  • This paper proposes MoDA, a novel multi-modal diffusion framework that generates high-fidelity talking head videos from a single portrait image, audio, and additional modalities.
@@
  • A coarse-to-fine fusion strategy is designed to progressively integrate noise motion with audio and other modalities, enabling effective cross-modal interaction.
@@ 3 Method
As illustrated in Fig. 3, MoDA is a two-stage framework. The first stage focuses on disentangling facial attributes, while the second stage generates motion based on audio and other conditions. The final output is rendered using the generated motion features and the 3D appearance features obtained from the first stage. For the first stage, we leverage the LivePortrait [13] framework to achieve fine-grained control over facial attributes, including expression, head pose, and gaze direction. Our primary contribution lies in the second stage, the motion feature generation network, which is built around four key components: Joint Parameter Space (see Sec. 3.2.1), Motion Transformer Blocks (see Sec. 3.2.2), Coarse-to-Fine Feature Fusion (see Sec. 3.2.3), and a carefully designed Loss Function for training (see Sec 3.2.4). These modules work collaboratively to ensure long-term identity consistency and natural facial expression generation.
3.1 Rectified Flow
@@
3.2 Motion Feature Generation with Diffusion Transformer
Similar to previous approaches [19, 44, 10], the primary conditioning signal for MoDA is the audio feature, incorporating additional auxiliary signals for the speaker’s identity and emotion to enhance controllability. However, instead of simply concatenating these conditioning signals, we propose to enhance the intrinsic identity and emotional information of the audio feature. This approach effectively balances identity and emotion across audio and speakers, enabling more natural lip control in real-world scenarios.
@@ 3.2.1 Joint Parameter Space
As shown in Fig. 3, we incorporate the existing facial reenactment framework [13] to extract disentangled facial representations. Specifically, the motion extractor $\mathcal{M}$ yields expression deformations $\delta$, head pose parameters $(R,t)$, the canonical keypoints of the source image $x_{c}$, and a scaling factor $s$. The motion representation $(R_{s},\delta_{s},t_{s},s_{s})\in\mathbb{R}^{70}$, serves as an identity-agnostic representation of the source input and is used to train MoDA to predict $(\hat{R},\hat{\delta},\hat{t},\hat{s})$ given audio input.
@@
Subsequently, the warping field estimator $\mathcal{W}$ computes a field from $x_{s}$ and $\hat{x}$ to deform the 3D features $f_{s}$, which are then passed to the generator $\mathcal{G}$ to synthesize the target image. The audio features $f_{a}$ are extracted using the wav2vec [29] encoder. To maintain consistency in the identity feature space across various scenarios, we use the canonical keypoints of the source image $x_{c}$ as identity information and generate identity features $f_{i}$. For facial emotion signals, we use a visual emotion classifier [28] to extract the speaker’s emotional labels and encode them into corresponding features $f_{e}$. The motion features from the first frame of each clip are used as guiding conditions to ensure inter-frame continuity and generate noisy motion features $f_{m}$. Inspired by EMO2 [37], emotion features $f_{e}$ are added to timestep embeddings to generate timestep features $f_{t}$, which are injected into each motion transformer block by adaptive layer normalization (AdaLN) [23]. AdaLN is used to prevent the degradation of emotion features during the joint-attention operation, ensuring that emotional cues are preserved throughout the interaction.
By integrating these conditioning signals, our model effectively generates realistic and temporally consistent motion features.
@@ 3.2.3 Coarse-to-Fined Feature Fusion
Although assigning modality-specific paths with separate QKV projections and FeedForward Networks (FFNs) in the attention mechanism can enhance cross-modal interaction, this design often overlook the inherent semantic commonalities shared across modalities. Independently learned weights hinder effective multimodal feature fusion and introduce redundant parameters, potentially causing convergence issues. This issue is particularly pronounced in tasks like audio-to-motion generation, where no pixel-level inputs are involved and the semantic gap between modalities is relatively small. In such cases, maintaining entirely separate parameterizations fails to provide meaningful benefits and instead leads to duplicated representations. To enhance the model’s capacity for multi-modal understanding, we introduce a Coarse-to-Fine Feature Fusion and Interaction strategy. This interaction strategy progressively integrates multi-modal features while reducing unnecessary parameters, thereby enhancing the model’s multi-modal understanding and training stability.
As illustrated in Fig. 3, the architecture progresses through three stages: a four-stream, two-stream, and single-stream process. In the four-stream stage, each modality is processed independently to learn modality-specific representations, facilitating early-stage feature differentiation and alignment. In the two-stream stage, the audio stream is concatenated with emotion and identity features to form a merged stream that shares weights. This design encourages a balanced integration of emotional and identity cues from both the audio and the reference image. Finally, all modalities are unified into a single representation stream to allow deeper cross-modal fusion and enhance the generative expressiveness of the model.
@@
Dataset Method FVD $\downarrow$ FID $\downarrow$ F-SIM $\uparrow$ Sync-C $\uparrow$ Sync-D $\downarrow$ Smooth (%) $\uparrow$
@@
Ours 214.182 42.001 0.913 5.757 8.346 0.9972
EchoMimic 0.870 2.292 12.130 48.657
JoyVASA 0.924 5.569 9.368 1.717
JoyHallo 0.825 7.000 8.167 59.735
Hallo 0.848 6.051 8.730 59.190
Hallo2 0.849 6.386 8.523 60.305
Ours (sync) 0.896 7.710 7.469 0.846
Ours 0.895 6.862 8.088 0.846
@@ 3.2.5 Classifier-free guidance (CFG)
3.2.6 Realtime Inference
The real-time conversational scenario is enabled by low-latency motion generation. We align audio features with the video frame rate and segment the audio into continuous 100-frame chunks for streaming generation. Additionally, leveraging low-dimensional intermediate representations and the rectified flow sampling approach, we reduce DiT inference denoising steps from 40 to 10 while maintaining comparable quality. For detailed comparison, please refer to the supplementary material.
@@ 4.1.1 Dataset and Metrics
Our motion generation model is primarily trained on three publicly available datasets: HDTF [52], CelebV-Text [49], and MEAD [41]. For evaluation, we conducted experiments on three distinct test sets. The first two are derived from public datasets, CelebV-HQ [53] and HDTF, each consisting of 50 randomly sampled clips ranging from 3 to 10 seconds in length. The third is an in-the-wild set with 20 diverse cases, including real individuals, animated characters, dynamic scenes, and complex headwear. Each sample is paired with audio that is speech, emotional dialogue, or singing. We utilize several evaluation metrics to assess the performance of the proposed method. The Fréchet Inception Distance (FID) [30] and the Fréchet Video Distance (FVD) [39] are used to assess the quality of the generated output, while the F-SIM [38] measures facial similarity. In addition, sync-C [4] and sync-D [4] metrics are introduced to evaluate lip synchronization between different methods. A temporal smoothness metric [16] is also utilized to monitor the continuity of generated motion.
@@ 4.2 Results and Analysis
We juxtapose the quantitative results of the proposed method against those of EchoMimic [3], JoyVASA [2], JoyHallo [31], Hallo [43], and Hallo2 [5]. We also compare two variants of our method: Ours (MoDA without ALSE) and Ours (Sync) (MoDA with ALSE).
@@
4.2.1 Quantitative Comparison
The quantitative results on the CelebV-HQ test and HDTF dataset are shown in Table 1. On the HDTF dataset, MoDA consistently outperforms all existing methods across all evaluation metrics. In particular, our method achieves the lowest FID and FVD scores, outperforming the second-best methods by 4.9 % and 16.0 %, respectively. This demonstrates the superiority of our method in terms of visual naturalness and the overall quality of the generated frames. The disentanglement of appearance and motion enables our method to better preserve identity consistency compared to existing approaches, as evidenced by the second-highest F-SIM score. Notably, JoyVASA achieves the highest F-SIM, likely due to its relatively constrained head movements, which contribute to higher frame-wise structural similarity.
On the CelebV-HQ test dataset, our method consistently outperforms the five baseline methods in all metrics except Sync-D, highlighting its robustness and effectiveness. Although MoDA slightly underperforms JoyHallo in synchronization confidence (Sync-C), it achieves a notable improvement in synchronization distance (Sync-D), along with further gains in motion smoothness. A detailed analysis can be found in Sec. 4.2.2. Table 1 also reports the results with and without incorporating ALSE. Introducing ALSE to supervise the synchronization between audio and keypoints leads to a significant improvement in lip-sync metrics.
Table 2 compares the performance of our method with existing approaches on the wild test dataset, which features more diverse characters and complex backgrounds. Our method consistently outperforms others on multiple metrics, demonstrating its robustness in unconfined scenarios. In addition, we evaluate the inference performance using the real-time factor (RTF). RTF < 1 indicates that the system meets the basic requirement for real-time inference. Our method achieves competitive efficiency, making it suitable for real-world applications that demand both quality and speed.

### v2

@@ Abstract
Talking head generation with arbitrary identities and speech audio remains a crucial problem in the realm of the virtual metaverse. Recently, diffusion models have become a popular generative technique in this field with their strong generation capabilities. However, several challenges remain for diffusion-based methods: 1) inefficient inference and visual artifacts caused by the implicit latent space of Variational Auto-Encoders (VAE), which complicates the diffusion process; 2) a lack of authentic facial expressions and head movements due to inadequate multi-modal information fusion. In this paper, MoDA handles these challenges by: 1) defining a joint parameter space that bridges motion generation and neural rendering, and leveraging flow matching to simplify diffusion learning; 2) introducing a multi-modal diffusion architecture to model the interaction among noisy motion, audio, and auxiliary conditions, enhancing overall facial expressiveness. In addition, a coarse-to-fine fusion strategy is employed to progressively integrate different modalities, ensuring effective feature fusion. Experimental results demonstrate that MoDA improves video diversity, realism, and efficiency, making it suitable for real-world applications. Project Page: https://lixinyyang.github.io/MoDA.github.io/
@@ 1 Introduction
Talking head generation creates a speaking portrait from a single image, guided by audio and other modalities. Despite advances in lip-synchronization, generating realistic and efficient videos remains a persistent challenge. The early methods [22, 25] based on Generative Adversarial Networks (GANs) [10] were the primary solution for audio-to-portrait translation, demonstrating relatively accurate lip-synchronization. However, these methods remain limited in capturing expressive emotions and natural head movements. Recently, diffusion models [7, 13, 31] have significantly advanced generative modeling, enabling the creation of highly diverse and high-fidelity images and videos. Motivated by this progress, significant research efforts [5, 3, 42, 2] have been made aimed at extending these innovations to the development of portrait image animation.
@@
These methods are typically categorized into single-stage and two-stage frameworks. The former [5, 36, 15, 41] rely on a VAE-style latent space to directly generate video content from the audio input, integrating both motion and appearance in one pass. Despite being conceptually simple and trainable from end to end, two major limitations still persist, as shown in Fig. 2: 1) inefficient inference process. Most diffusion-based approaches rely on VAEs to model both lip motion diversity and visual appearance, but the implicit latent space and decoding complexity of VAEs significantly increase computational overhead, limiting their suitability for real-time applications; 2) limited capability for long-term video generation and visual artifacts. Methods like EMO [36] generate new video content by conditioning on previous frames. These prior frames convey motion cues but also leak appearance details, causing pixel errors to accumulate and degrade video quality.
The two-stage methods [42, 2, 17] use diffusion models to generate low-dimensional intermediate motion representations, simplifying the learning process. However, the accuracy of these intermediate representations, and therefore the final video quality, is often limited by inadequate multi-modal fusion. A key reason is that conventional cross-attention allows information to flow in just one direction, typically from audio to motion, preventing audio cues from being refined with motion context. As shown in ablation study, this limitation produces motion sequences that become inconsistent when the model is conditioned on arbitrary identities inputs.
@@
In this paper, we propose a novel framework named MoDA, designed to generate realistic talking head videos. MoDA first establishes a joint parameter space that bridges motion generation and neural rendering, consisting of a disentangled motion-appearance space [12] along with the corresponding audio, emotion, and identity information. To address the fusion bottleneck, we draw inspiration from recent lip-to-speech tasks [38, 24], visual information can provide contextual cues complementary to audio. As shown in Fig. 3, MoDA introduces the Multi-modal Diffusion Transformer (MMDiT) [9], equipped with rectified flow [19], as the backbone diffusion framework to facilitate multi-modal fusion. In this design, MoDA can dynamically adapt audio features based on motion, identity, and emotion, thereby improving the accuracy of motion generation.
@@
Another cause of inadequate multi-modal fusion is the frequent neglect of semantic overlaps within the joint parameter space: for example, audio inherently contains emotional and identity-related cues. Independently learned weights prevent the model from fully exploiting these intrinsic correlations, ultimately hindering its overall performance. Therefore, a coarse-to-fine fusion architecture is further proposed to progressively enhance feature integration across modalities. Initially, the model maintains modality-specific weights to capture the unique characteristics of each condition. In the intermediate stage, audio and auxiliary conditions, such as identity and emotion, begin to share attention weights. Finally, all modalities are merged into a unified representation space, allowing holistic fusion and improving the overall generative expressiveness. Moreover, to further encourage precise lip-synchronization while maintaining motion diversity, this paper provides an optional Adaptive Lip-motion Synchronization Expert (ALSE), which can be integrated during training. The contributions of MoDA can be summarized as follows:
@@
  • This paper proposes MoDA, a novel multi-modal diffusion framework that generates high-fidelity talking head videos from an image, audio, and additional modalities.
@@
  • A coarse-to-fine fusion strategy is designed to progressively integrate noisy motion with audio and other modalities, enabling effective multi-modal fusion.
@@ 3 Method
3.1 Preliminaries
@@
3.2 Model Architecture
Instead of simply concatenating these conditioning signals, we enhance the intrinsic characteristics and emotional nuances of the audio representation by treating external emotion and identity cues as "catalysts". This approach balances identity and emotion between audio and speakers, enabling more natural lip control in real-world scenarios.
@@ 3.2.1 Joint Parameter Space
As shown in Fig. 3, we incorporate the existing facial re-enactment framework [12] to extract disentangled facial representations. Specifically, the motion extractor $\mathcal{M}$ yields expression deformations $\delta$, head pose parameters $(R,t)$, the canonical keypoints of the source image $x_{c}$, and a scaling factor $s$. The motion representation $(R_{s},\delta_{s},t_{s},s_{s})\in\mathbb{R}^{70}$, serves as an identity-agnostic representation of the source input and is used to train MoDA to predict $(\hat{R},\hat{\delta},\hat{t},\hat{s})$ given audio input.
@@
Subsequently, the warping field estimator $\mathcal{W}$ computes a field from $x_{s}$ and $\hat{x}$ to deform the 3D features $f_{s}$, which are then passed to the generator $\mathcal{G}$ to synthesize the target image. The audio features $f_{a}$ are extracted using the wav2vec [27] encoder. To maintain consistency in the identity feature space across various scenarios, we use the canonical keypoints of the source image $x_{c}$ as identity information and generate identity features $f_{i}$. For facial emotion signals, we use a visual emotion classifier [26] to extract the speaker’s emotional labels and encode them into corresponding features $f_{e}$. The motion features from the first frame of each clip are used as guiding conditions to ensure inter-frame continuity and generate noisy motion features $f_{m}$. Inspired by EMO2 [35], emotion features $f_{e}$ are added to timestep embeddings to generate timestep features $f_{t}$, which are injected into each motion transformer block by adaptive layer normalization (AdaLN) [21]. AdaLN is used to prevent the degradation of emotion features during the joint-attention operation, ensuring that emotional cues are preserved throughout the fusion. By integrating these conditioning signals, our model effectively generates realistic and temporally consistent motion features.
@@ 3.2.3 Coarse-to-Fined Feature Fusion
Although assigning modality-specific paths with separate QKV projections and FeedForward Networks (FFNs) in the attention mechanism can enhance multi-modal fusion, this design often overlook the inherent semantic commonalities shared across modalities. Independently learned weights hinder effective multi-modal feature fusion and introduce redundant parameters, potentially causing convergence issues. This issue is particularly pronounced in tasks like audio-to-motion generation, where no pixel-level inputs are involved and the semantic gap between modalities is relatively small. In such cases, maintaining entirely separate parameterizations fails to provide meaningful benefits and instead leads to duplicated representations. To enhance the model’s capacity for multi-modal understanding, we introduce a Coarse-to-Fine Feature Fusion strategy. This fusion strategy progressively integrates multi-modal features while reducing unnecessary parameters, thereby enhancing the multi-modal understanding and training stability.
As illustrated in Fig. 3, the architecture progresses through three stages: a four-stream, two-stream, and single-stream process. In the four-stream stage, each modality is processed independently to learn modality-specific representations, facilitating early-stage feature differentiation and alignment. In the two-stream stage, the audio stream is concatenated with emotion and identity features to form a merged stream that shares weights. This design encourages a balanced integration of emotional and identity cues from both the audio and the reference image. Finally, all modalities are unified into a single representation stream to allow deeper fusion and enhance the generative expressiveness.
@@
Dataset Method FVD $\downarrow$ FID $\downarrow$ F-SIM $\uparrow$ Sync-C $\uparrow$ Sync-D $\downarrow$ Smo(%) $\uparrow$
@@
Ours 205.442 44.071 0.913 5.878 8.135 0.9972
@@ 3.2.4 Loss Function
3.3 Realtime Inference
The real-time conversational scenario is enabled by low-latency motion generation. We align audio features with the video frame rate and segment the audio into continuous 100-frame chunks for streaming generation. Additionally, leveraging low-dimensional intermediate representations and the rectified flow, we reduce DiT inference denoising steps from 50 to 10, yet achieve even higher quality. For detailed comparison, please refer to the supplementary material.
@@ 4.1.1 Dataset and Metrics
Our motion generation model is primarily trained on three publicly available datasets: HDTF [49], CelebV-Text [46], and MEAD [39]. For evaluation, we conducted experiments on three distinct test sets. The first two are derived from public datasets, CelebV-HQ [50] and HDTF, each consisting of 50 randomly sampled clips ranging from 3 to 10 seconds in length. The third is an in-the-wild set with 20 diverse cases, including real individuals, animated characters, dynamic scenes, and complex headwear. Each sample is paired with audio that is speech, emotional dialogue, or singing. We utilize several evaluation metrics to assess the performance of the proposed method. The Fréchet Inception Distance (FID) [28] and the Fréchet Video Distance (FVD) [37] are used to assess the quality of the generated output, while the F-SIM [36] measures facial similarity. In addition, Sync-C [4] and Sync-D [4] metrics are introduced to evaluate lip-synchronization between different methods. A temporal smoothness metric (Smo) [14] is also utilized to monitor the continuity of generated motion.
@@ 4.1.2 Implementation Details
EchoMimic 0.870 2.292 12.130 48.657
JoyHallo 0.825 7.000 8.167 59.735
Hallo 0.848 6.051 8.730 59.190
Hallo2 0.849 6.386 8.523 60.305
Ditto 0.923 6.107 9.040 0.792
JoyVASA 0.924 5.569 9.368 1.717
Ours (sync) 0.896 7.710 7.469 0.846
Ours 0.895 6.862 8.088 0.846
@@ 4.2 Results and Analysis
We juxtapose the results of the proposed method against those of EchoMimic [3], Ditto [17], JoyVASA [2], JoyHallo [29], Hallo [41], and Hallo2 [5], Ours and Ours (Sync) (MoDA with ALSE).
4.2.1 Quantitative Comparison
The quantitative results on the CelebV-HQ test and HDTF dataset are shown in Table 1. On the HDTF dataset, MoDA consistently outperforms all existing methods across all evaluation metrics. In particular, our method achieves the lowest FID and FVD scores, outperforming the second-best methods by 4.9 % and 16.0 %, respectively. This demonstrates the superiority of our method in terms of visual naturalness and the overall quality of the generated frames. The two-stage methods better preserve identity, as evidenced by their F-SIM scores. Notably, JoyVASA and Ditto achieve the highest and second-highest F-SIM, respectively—likely due to their relatively constrained head movements and facial expressions, which enhance frame-wise structural similarity. The correspondingly lower FVD scores further reflect these limitations. In contrast, MoDA effectively mitigates such issues, demonstrating that enhancing multi-modal fusion can lead to improved overall model performance.
On the CelebV-HQ test dataset, MoDA consistently outperforms the six baseline methods in all metrics except Sync-D and Smo, highlighting its robustness. Although MoDA slightly underperforms JoyHallo in synchronization confidence (Sync-C), it achieves a notable improvement in synchronization distance (Sync-D), along with further gains in motion smoothness. Tables 1 also reports the results with and without ALSE. Introducing ALSE to supervise the synchronization between audio and keypoints leads to a significant improvement in lip-sync.
Table 2 shows that our method outperforms existing approaches across multiple metrics on the wild test set with diverse identities. We also report the real-time factor (RTF), where RTF <1 indicates real-time capability. Ditto achieves the lowest RTF, benefiting from TensorRT acceleration. Our method also delivers competitive efficiency, demonstrating its suitability for real-world applications.